### v1

# AnimateScene: Camera-controllable Animation in Any Scene

Abstract
3D scene reconstruction and 4D human animation have seen rapid progress and broad adoption in recent years. However, seamlessly integrating reconstructed scenes with 4D human animation and produce visually engaging results remains challenging. One key difficulty lies in placing the human at the correct location and scale within the scene while avoiding unrealistic interpenetration. Another challenge is that the human and the background may exhibit different lighting and style, leading to unrealistic composites. In addition, appealing character motion videos are often accompanied by camera movements, which means that the viewpoints need to be reconstructed along a specified trajectory. We present AnimateScene, which addresses above issues in a unified framework. First, we design an accurate placement module that automatically determines a plausible 3D position for the human and prevents any interpenetration within the scene during motion. Second, we propose a training‑free style alignment method that adapts the 4D human representation to match the background’s lighting and style, achieving coherent visual integration. Finally, we design a joint post‑reconstruction method for both the 4D human and the 3D scene that allows camera trajectories to be inserted, enabling the final rendered video to feature visually appealing camera movements. Extensive experiments show that our AnimateScene generates dynamic scene videos with high geometric detail and spatiotemporal coherence across various camera and action combinations.
Figure 1: Camera-controllable 4D human animation with AnimateScene. AnimateScene takes a single human source image and a single-view scene image as input. It reconstructs both an animatable 3D Gaussian human and a 3D Gaussian scene, and then outputs a dynamic rendered video. In this video, the camera follows a user-defined trajectory while the actor performs the driving motion. This yields a style-consistent, collision-free integration of human motion and environment.
1 Introduction
Camera-controllable animation can build dynamic scenes that preserve spatiotemporal consistency with both subject motion and viewpoint shifts, yielding more realistic scene composition. Recent studies have explored injecting camera (Liu et al. 2024b, a; Zhu et al. 2024; Zeller 2024; Huang et al. 2025) or human motion control (Liu et al. 2021a; Gao et al. 2023) in the 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) framework, while there is no unified camera-and-human control solution for 3DGS. This task takes a single 2D scene image, a 2D human image, and a target camera trajectory as motion control inputs, result in a synthesized scenes that can provide vivid camera-controllable backgrounds for 4D human sequence. Some recent studies explore it in video generation (Cao et al. 2025; Wang et al. 2024d). Despite these advances, video generation models face significant limitations (Wang et al. 2025), including lengthy inference time and lack of 3D constraints.
To generate dynamic videos with controllable camera and human motion from a single scene image and a single human image with motion, a trivial method is to directly combine single-view 3D scene reconstruction (Che et al. 2024; Zhao et al. 2024; Yu et al. 2024a) with animatable 3D motion reconstruction (Yu et al. 2024b; Hou et al. 2025). However, there are several challenges in seamlessly integrating the character with the scene and rendering realistic and coherent camera motion in the generated video. A major challenge is accurately positioning and scaling the person in the scene without causing implausible overlap. Another challenge arises when foreground and background differ in lighting and style, resulting in unconvincing composites. Additionally, engaging character-motion videos often involve moving cameras, requiring the reconstruction of viewpoints along a defined trajectory.
In this paper, we propose AnimateScene to address the above challenges. For the first challenge, AnimateScene tackles this placement challenge with a depth-guided 3D module. Specially, we pool depth values along the bounding-box baseline and back-project the bottom-centre pixel to world space and rescale the avatar so that its projection matches the 2D box size. If the avatar collides with the scene, we move it to the nearest free space. At the same time, a light temporal smoother keeps the actor collision-free throughout the entire motion clip. For the second challenge, we employ a training-free style transfer module that injects the background’s lighting and color palette into the foreground while leaving its geometry intact. This produces a style-consistent avatar ready for seamless compositing. For the third challenge, we propose joint post-reconstruction to render the fused human–scene Gaussian field along the planned camera path. The joint post-reconstruction module detects any newly exposed gaps, dilates the masks, and fills them with diffusion-based inpainting. Because this happens after the avatar is inserted, it eliminates leftover foreground-background collisions and produces smoother, more consistent composites in every moving-camera view.
In summary, AnimateScene begins with four straightforward inputs: a single scene image, a single human image, an accompanying motion clip, and a user-defined camera trajectory. AnimateScene first applies a training-free style-transfer module that aligns the human image to the scene’s lighting and colour palette while preserving geometry. With appearance harmonised, the person is reconstructed as an animatable 4D Gaussian avatar sequence, and a single-view reconstructor generates a sparse 3D Gaussian background. A depth-guided placement network subsequently converts the 2D bounding box into a collision-free 3D location and scale, positioning the moving actor physically correctly within the scene. AnimateScene concludes with a joint post-reconstruction stage that refines the fused human–scene field along the entire camera path, inpainting any newly exposed regions to yield smooth, artefact-free renders from every viewpoint.
Across extensive experiments on a diverse set of indoor and outdoor environments with varying camera trajectories and human motion sequences, AnimateScene consistently surpasses state-of-the-art baselines in rendering quality, geometric consistency, and placement plausibility, validating its robust generalization to dynamic 4D scenarios. In summary, our main contributions are as follows.
  • We introduce AnimateScene, a unified framework that couples single-image scene reconstruction, style-aligned 4D humans and controllable camera paths, enabling joint control over background and actor motion.
  • We propose a depth-guided 3D placement module that upgrades 2D object-placement predictions into collision-free world coordinates, guaranteeing physically coherent integration of animatable 3D Gaussian avatars.
  • We propose a joint post-reconstruction module to eliminate foreground-background occlusion, and ensure geometric and stylistic consistency across varied camera–action combinations.
  • Extensive experiments demonstrate that AnimateScene achieves rendering with rich geometric detail and spatiotemporal coherence across diverse camera trajectories and human action combinations, significantly advancing dynamic scene generation.
2 Related work
3D scene generation. Over the past decade, the field of 3D scene generation has undergone a profound transformation, moving from manually designed rule based pipelines (Mandelbrot 1967; Fournier, Fussell, and Carpenter 1982; Müller et al. 2006; Parish and Müller 2001; Cordonnier et al. 2023; Paris et al. 2023; Vanegas et al. 2012; Emilien et al. 2015) toward fully learned photorealistic synthesis (Wang et al. 2018; Bahmani et al. 2023; Lin et al. 2023; Po and Wetzstein 2024; Cohen-Bar et al. 2023; Zhou et al. 2024; Chung et al. 2023; Yu et al. 2024b; Hou et al. 2025; Che et al. 2024; Zhao et al. 2024; Yu et al. 2024a; Wu et al. 2025a). With the advent of deep generative models, researchers began to train Generative Adversarial Networks (GANs) (Wang et al. 2018; Bahmani et al. 2023; Lin et al. 2023) and more recently diffusion networks to directly model complex spatial distributions of real world environments, when coupled with powerful implicit representations such as Neural Radiance Fields (NeRFs) (Po and Wetzstein 2024; Cohen-Bar et al. 2023; Zhou et al. 2024) or 3D Gaussian splatting (Chung et al. 2023; Yu et al. 2024b; Hou et al. 2025), these neural 3D approaches achieved unprecedented fidelity, view consistency, and rendering efficiency. Parallel strands of work (Yu et al. 2024c; Yang et al. 2025; Wang et al. 2024a; Li et al. 2023; Tan et al. 2024) recast scene generation as an image or video synthesis problem either by holistically outpainting panoramas or iteratively extending views, thereby leveraging advances in text or image conditioned diffusion to produce large scale, diverse environments, albeit often at the expense of precise geometric depth. The latest video based paradigms (Che et al. 2024; Zhao et al. 2024; Yu et al. 2024a; Wu et al. 2025a) further marry temporal coherence with spatial structure by treating scene creation as a dynamic video generation task, employing state of the art video diffusion models to synthesize fully navigable, time varying worlds.
Motion-Conditioned Video Generation. With the rapid advancement of diffusion models (Peebles and Xie 2023; Ramesh et al. 2022; Rombach et al. 2022), various efforts have been made to enhance the performance of video generation, including improvements in model architecture (Jin et al. 2025; Ma et al. 2024), learning strategies (Yang et al. 2024; Singer et al. 2023), and data curation (Qiu et al. 2023; Wang et al. 2023, 2024c). Recently, motion-conditioned video generation has witnessed significant advancements (Wu et al. 2024, 2025b; Gao et al. 2025; Yu et al. 2024b; Hou et al. 2025). Recently, motion-conditioned video generation has witnessed significant advancements (Wu et al. 2024, 2025b; Gao et al. 2025; Yu et al. 2024b; Hou et al. 2025). MotionBooth (Wu et al. 2024) introduces a two-stage method—an initial fine-tuning phase followed by training-free refinement—that combines subject-token cross-attention manipulation with a latent-shift module for precise subject and camera motion control. MotionMatcher (Wu et al. 2025b) refines text-to-video diffusion models by matching spatio-temporal motion features, enabling precise, leak-free motion control. MotionDistill (Janson, Popa, and Belilovsky 2024) introduces a differentiable pipeline to distill fine-grained motion from pretrained models. TokenMotion (Li et al. 2025) applies token disentanglement in DiT-based diffusion to separately steer camera and pose. SynFMC (Shuai et al. 2025) provides a fully annotated synthetic dataset for benchmarking free-form object and camera motion control. Identity-preserving frameworks like MotionCharacter (Fang et al. 2024) add modules to maintain subject identity during nuanced edits. Guided Motion Diffusion (Karunratanakul et al. 2023) embeds spatial constraints and trajectory priors into denoising to follow predefined paths. MotionCtrl (Wang et al. 2024d) proposes a unified motion controller that decouples camera and object motion, allowing independent and flexible trajectory control in video diffusion frameworks. Different from the above methods, our main task is to generate dynamic videos with controllable camera and human motion from a single scene image and human image.
Object placement. Methods fall into generative (Zhou et al. 2025a; Qin et al. 2024; Wu et al. 2023; Hong et al. 2024; Zhang et al. 2023; Zhou et al. 2022; Tripathi et al. 2019; Lin et al. 2018), which propose transformations for realism, and discriminative (Gao, Zhang, and Niu 2025; Zhu et al. 2023; Niu et al. 2022; Dvornik, Mairal, and Schmid 2019; Yuan et al. 2023; Liu et al. 2021b), which score or regress placements. Generative examples include ST-GAN (Lin et al. 2018), which iteratively refines Spatial Transformer warps in a GAN, and SF-GAN (Dvornik, Mairal, and Schmid 2019), which jointly optimizes geometry and appearance. Discriminative approaches like FOPA (Niu et al. 2022) predict full-score maps using dynamic filters and background priors, while TopNet (Zhu et al. 2023) learns the correlation between foregrounds and backgrounds, producing a 3D heatmap containing the rationality scores of all locations and scales. Prior object placement methods drives realistic 2D placement, either by learning warp transformations or by scoring/regressing boxesm, but it remains bound to the image plane and ignores depth.
3 Method
As illustrated in Fig. 2, AnimateScene consists of three main parts, 1) an animatable human style transfer module, 2) an object placement model, as well as 3) a trajectory‑based joint human–scene reconstruction model. Given an input RGB human image $I_{h}\in\mathbb{R}^{H\times W\times 3}$, an input RGB scene image $I_{s}\in\mathbb{R}^{H\times W\times 3}$, a target camera trajectory $\boldsymbol{T}=\left\{\boldsymbol{T}_{\mathrm{i}}\right\}_{\mathrm{i}=1}^{\mathrm{n}}\in\mathbb{R}^{\mathrm{n}\times 4\times 4}$, and a human motion video $V_{h}\in\mathbb{R}^{F\times H\times W\times 3}$, we first use the style transfer module to align the style of the human image $I_{h}$ with that of the scene image $I_{s}$. Then the animatable human reconstruction model can reconstruct an style-consistent animatable 4D human avatar sequence. Single-view scene reconstruction model is utilized to reconstruct a 3D scene from a single-view image $I_{s}$. The object placement model predicts a reasonable placement for $I_{h}$ in terms of $I_{s}$ , involving location $(x,y,z)$ and scale $(w,h)$. The animated 3D gaussian human is placed in the reconstructed 3D scene with the guidance of the predicted placement. Finally, we jointly reconstruct the fused 3D scene along the camera trajectory $\boldsymbol{T}$.
Figure 2: Overview of AnimateScene. AnimateScene takes a single scene image, a single human image, an accompanying motion clip, and a user-defined camera path. It first aligns the human’s appearance with the scene. Next, it reconstructs a 4D Gaussian avatar alongside a sparse 3D Gaussian scene. Depth cues are then used to lift the avatar to a collision-free 3D location and scale. The system finally refines the fused human–scene field along the camera trajectory, inpainting any newly exposed regions. The result is a video where actor, environment, and viewpoint move in seamless geometric and stylistic harmony.
3.1 Preliminaries
3D Gaussian Splatting. 3D Gaussian Splatting (Kerbl et al. 2023) represents a scene as a collection of anisotropic Gaussian primitives, each defined by a centroid $\mathbf{x}\in\mathbb{R}^{3}$, scale parameters $\mathbf{s}\in\mathbb{R}^{3}$, and a rotation quaternion $\mathbf{q}\in\mathbb{R}^{4}$, together with an opacity $\alpha\in\mathbb{R}$ and an appearance feature vector $\mathbf{c}\in\mathbb{R}^{C}$. View-dependent shading is captured by encoding $\mathbf{c}$ with spherical harmonics. During rendering, each 3D Gaussian is projected into the image plane as a 2D Gaussian and composited front-to-back via alpha blending.
Animatable Human Reconstruction. We adopt the SMPL/SMPL‑X (Loper et al. 2015; Pavlakos et al. 2019) parametric body models, which represent the human mesh using shape parameters $\boldsymbol{\beta}\in\mathbb{R}^{20}$ and pose parameters $\boldsymbol{\theta}\in\mathbb{R}^{55\times 3}$, enabling low‑dimensional, deformable articulation. For animatable human reconstruction from a single RGB image, we employ the Large Animatable Human Model (LHM) (Qiu et al. 2025), which generalizes to in‑the‑wild inputs and supports real‑time, pose‑controllable avatar creation. Given a human image (optionally with SMPL‑X pose or motion sequence), LHM extracts SMPL‑X anchors and 2D visual tokens via a frozen vision encoder, fuses them in a multimodal transformer, and decodes per‑Gaussian parameters (position, anisotropic scale, rotation, opacity, spherical‑harmonic appearance). The resulting canonical 3D Gaussian representation preserves facial identity, clothing geometry, and texture, and can be skinned via linear blend skinning for photo‑realistic rendering and seamless animation.
Single-View Scene Reconstruction. We utilize VistaDream (Wang et al. 2024b) to reconstructs our 3D scene, which completes 3D scene from a single input image by progressively expanding and refining its content. It first expands the field of view through synthetic zoom-out and fills missing areas using diffusion-based inpainting guided by scene-specific text prompts, producing a complete RGB image and a monocular depth map to form a coarse 3D Gaussian scaffold. The method then warps this scaffold to novel viewpoints along the target camera trajectory, inpaints unseen regions, re-estimates depth, and incrementally augments the Gaussian field, improving coverage with each iteration. To achieve strict view-to-view consistency, VistaDream renders multiple viewpoints from the evolving Gaussian scene and applies view fusion inpainting, where diffusion-based denoising is blended with scaffold renderings to enforce cross-view coherence. The refined multi-view results are finally used to fine-tune the Gaussian field, yielding a high-fidelity, geometrically aligned 3D scene renderable from arbitrary novel viewpoints.
3.2 Style Transfer Module
We employ Stable Diffusion (Rombach et al. 2022) equipped with IP‑Adapter (Ye et al. 2023) to align the appearance of the human image $I_{h}$ with the style of the background scene image $I_{s}$. Specifically, the style reference image $I_{s}$ is fed into the IP-Adapter branch of Stable Diffusion, while the subject image $I_{h}$ is provided as the primary input for stylization. The IP‑Adapter modulates the diffusion process to inject the background’s style into the human rendering without altering its geometry or pose. This process produces a style-consistent human representation that can be seamlessly integrated into the reconstructed scene. Then we feed the style transfered and the driven video $V_{h}$ into the the Large Animatable Human Model (LHM) (Qiu et al. 2025) to generate the 4D human sequence.
With style-consistent human and scene images, we reconstruct the 4D human sequence and the 3D scene using LHM (Qiu et al. 2025) and VistaDream (Wang et al. 2024b), respectively. Specifically, we feed the human image $I_{h}$ and driving video $V_{h}$ into LHM to obtain a 4D human motion sequence that follows the motion in the driving video. Unlike VistaDream, which directly reconstructs the entire scene along a predefined camera trajectory, we only use the depth estimation part of VistaDream and obtain a sparse 3D representation by projecting scene pixels into 3D space. This enables a more accurate estimation of the 4D human sequence’s position within the scene. Once the human position is determined, performing inpainting and reconstruction along the camera trajectory prevents overlap between the 4D human Gaussians and the 3D scene Gaussians.
3.3 Object Placement Module
Existing object placement methods operate exclusively in the two‐dimensional image domain, where a network predicts a bounding box $\mathbf{B}=(x,y,w,h)$, where $(x,y)$ denotes the top‐left corner and $(w,h)$ its width and height in the image $I_{h}$. Our task, however, requires estimating a plausible three‐dimensional location for an animatable 4D Gaussian human within a reconstructed 3D Gaussian scene, a problem that cannot be directly solved by these 2D‐focused models.
The set of $N$ uniformly spaced points along the bottom edge of this box is denoted by $\mathcal{E}=\{(u_{n},v_{n})\}_{n=1}^{N}$, and the depth map of the reconstructed Gaussian scene is given by $D(u,v)$. The camera intrinsic matrix is defined as
$K=\begin{bmatrix}f_{x}&0&c_{x}\\ 0&f_{y}&c_{y}\\ 0&0&1\end{bmatrix}$, (1)
and the resulting three‐dimensional insertion point is $\mathbf{p}^{*}=(X^{*},Y^{*},Z^{*})$ in world coordinates.
First, an off‐the‐shelf 2D placement network is applied to predict the bounding box $\mathbf{B}=(x,y,w,h)$ for the human in the image $I$. Next, we sample $N$ points along the bottom edge of this box by setting
$u_{n}=x+\frac{n}{N}\,w,\quad v_{n}=y+h,\quad n=1,\dots,N,$ (2)
thereby defining $\mathcal{E}$. To estimate the depth coordinate $Z^{*}$, we compute the average of the depth map values at these sampled points,
$Z^{*}=\frac{1}{N}\sum_{n=1}^{N}D(u_{n},v_{n}),$ (3)
which mitigates noise from single‐pixel errors by aggregating information across the box baseline. We then select the center pixel of the bottom edge,
$(u_{c},v_{c})=\Bigl{(}x+\tfrac{w}{2},\,y+h\Bigr{)},$ (4)
and back‐project it into 3D space using the inverse intrinsics:
$\mathbf{p}^{*}=Z^{*}\,K^{-1}\begin{bmatrix}u_{c}\\ v_{c}\\ 1\end{bmatrix}=\begin{bmatrix}X^{*}\\ Y^{*}\\ Z^{*}\end{bmatrix}$. (5)
Finally, the 4D Gaussian human is positioned at $\mathbf{p}^{*}$ with a spatial covariance chosen so that its projection under the camera model matches the original 2D size $(w,h)$. To avoid penetration of the 4D Gaussian human into the reconstructed 3D Gaussian scene during motion, we enforce a collision-free constraint in world coordinates. Let $\mathcal{I}(\mathbf{q})$ be an occupancy indicator of the background Gaussian field, where $\mathcal{I}(\mathbf{q})=1$ denotes that $\mathbf{q}$ lies in a high-density background region, and $\mathcal{I}(\mathbf{q})=0$ otherwise. If $\mathcal{I}(\mathbf{p}^{*})=1$, we project $\mathbf{p}^{*}$ to the nearest free-space point:
$\mathbf{p}^{*}\leftarrow\arg\min_{\mathbf{q}\in\mathbb{R}^{3},\ \mathcal{I}(\mathbf{q})=0}\|\mathbf{q}-\mathbf{p}^{*}\|^{2}.$ (6)
Even with collision-free projection, small frame-to-frame jitters or local overlaps may occur, so we apply temporal smoothing to the sequence of insertion points $\{\mathbf{p}^{*}_{t}\}_{t=1}^{T}$:
$\mathbf{p}^{*}_{t}\leftarrow(1-\lambda)\,\mathbf{p}^{*}_{t}+\lambda\,\frac{\mathbf{p}^{*}_{t-1}+\mathbf{p}^{*}_{t+1}}{2},$ (7)
where $\lambda\in[0,1]$ controls the smoothing strength. This step guarantees a natural, continuous trajectory without abrupt spatial shifts while maintaining collision-free placement.
3.4 Joint Post‑Reconstruction
Once the 4D human motion sequence is properly inserted into the 3D scene, we follow the VistaDream(Wang et al. 2024b) approach to reconstruct the fused scene along the predefined camera trajectory. To obtain higher-quality videos, after merging the reconstructed 3D scene with the 4D human and rendering along the target camera trajectory, we further performed post-processing on the video. Specifically, we selected the background color that contrasts significantly with the synthesized scene and we automatically detect all remaining unfilled cavities visible in the reconstructed color volume, resulting a binary mask that isolates these hole regions. To ensure seamless coverage of their boundaries, this mask is then expanded via a small radius morphological dilation. The dilated mask and the original reconstruction are passed together through our learned inpainting network, which hallucinates plausible content for each masked region by leveraging the surrounding geometry and appearance cues. Finally, the newly inpainted output is composited back into the 3D model, yielding a reconstruction with substantially reduced artifacts and holes while maintaining geometric fidelity and photorealistic detail.
Figure 3: Comparison results of our method and baselines. Each row shows the scene image (col. 1), the human (col. 2), the synthesized result of 3DitScene (col. 3), the synthesized result of SEVA (col. 4), the synthesized result of DimensionX (col. 5) and the synthesized result of our method (col. 6). It can be seen that our method achieves the best results in both image quality and geometric consistency when the camera rotates to unseen viewpoint.
4 Experiments
4.1 Experimental setup
Datasets. We utilize 83 single-view, copyright-free online photos and 83 generated 4D Human animation samples for evaluation. The image set encompasses both indoor and outdoor environments, spanning real-world and simulated scenes. The 4D Human data features a diverse range of dance and sports movements generated by LHM (Qiu et al. 2025). Each 4D Human instance is paired with a carefully selected scene to best complement its motion characteristics. We designed six types of trajectories to evaluate the quality of scene generation with controllable cameras. More details can be found in the supplementary material.
Evaluation metrics. Drawing inspiration from previous works (Wang et al. 2024b; Wang, Chan, and Loy 2023), we adopt the Vision-Language Model LLaVA (Liu et al. 2023) to evaluate the rendered synthetic scene video frames with different camera trajectories. The LLaVA-IQA metric focuses on five dimensions: noise suppression, sharpness of edges, structural consistency, level of detail, and overall perceptual quality. Details of the LLaVA-IQA metric are given in supplementary.
Baselines. Since there is currently no method for reconstructing camera controllable 3D background scenes for 4D humans, we compare our method with one 3D scene reconstruction work 3DitScene (Zhang et al. 2024) and two generative view synthesis method SEVA (Zhou et al. 2025b) and DimensionX (Sun et al. 2024). 3DitScene is a unified scene editing framework leveraging language-guided disentangled gaussian splatting that enables seamless editing from 2D to 3D, allowing generate 3D scene from single image and synthesize novel viewpoints. SEVA and DimensionX are novel view video generation method. SEVA is a generalist diffusion model that creates novel views of a scene, given any number of input views and target cameras. DimensionX is a framework designed to generate photorealistic 3D and 4D scenes from just a single image with video diffusion.
4.2 Evaluation results
Method Noise-Free $\uparrow$ Edge $\uparrow$ Sturcture $\uparrow$ Detail $\uparrow$ Quality $\uparrow$
3DitScene (Zhang et al. 2024) 0.696 0.262 0.533 0.645 0.538
SEVA (Zhou et al. 2025b) 0.432 0.045 0.291 0.429 0.312
DimensionX (Sun et al. 2024) 0.759 0.135 0.392 0.817 0.559
Ours $\mathbf{0.980}$ $\mathbf{0.344}$ $\mathbf{0.636}$ $\mathbf{0.979}$ $\mathbf{0.759}$
Table 1: LLaVA-IQA scores on synthetic video frames. AnimateScene ranks highest on five metrics (noise suppression, edge sharpness, structural consistency, detail level, and overall perceptual quality).
3DitScene (Zhang et al. 2024) SEVA (Zhou et al. 2025b) DimensionX (Sun et al. 2024) Ours
Consistency Human 17.26 7.84 18.91 55.99
GPT4-v 11.37 4.68 22.05 61.90
Quality Human 20.42 3.77 24.31 51.50
GPT4-v 17.68 6.93 28.47 46.92
Table 2: The percentage of favorite users for the video quality and geometric consistency of video generated by each method. Human and GPT-4v vote shares for each method’s video quality and geometric consistency, with AnimateScene leading both.
Quantitative results. We provide a project page (HTML) in the supplementary materials, a local HTML file (project_page.html), which exhibits the videos generated by our method and baselines. To view the qualitative results, please open project_page/project_page.html in a web browser. Since there is currently no method for reconstructing camera controllable 3D background scenes for 4D humans, we composite the baseline-generated controllable-view 3D scenes with the 4D human sequences frame by frame. Specifically, we render the human sequences according to the predefined camera trajectory, and then paste each rendered human frame onto the baseline-generated background according to positions provided by the object placement model, which ensures geometric consistency between the human and the scene in the final video. In Table 1, we report the LLaVA scores across different dimensions, which showed that our method is capable of synthesizing high-quality scenes with controllable camera trajectories.
To further evaluate the synthetic result, we conduct a user study to compare the rendered video by our method with the established baselines. We ask 20 users to vote for their preferred method based on video quality and geometric consistency across different viewpoints. Followed 3DitScene (Zhang et al. 2024), we also adopt GPT-4v as an additional criterion since it can evaluate 3D consistency and image quality well aligned with human preference. In Table 2, we report the user study result, which shows our method outperforms other baselines in terms of both video quality and geometric consistency.
Figure 4: Ablation study on object placement and scene reconstruction. Each row shows the scene (col. 1), human (col. 2), baseline without key components (col. 3), and our full AnimateScene (col. 4). In rows 1–2, 2D placement yields a sinking foot or floating figure, whereas ours produces stable, natural poses. In rows 3–4, skipping view‐fusion inpainting and post‐processing leaves black holes by the sofa and road; our full method fills them for a coherent 3D scene.
Qualitative results. In Figure 3, we show qualitative results of our methods and other baselines. Since different reconstruction can render different view under the same camera position, and the positional information is encoded as a condition in video generation models, the corresponding frames produced by each method still differ slightly even we use the same camera trajectory. It can be seen that 3DitScene (Zhang et al. 2024) produces noticeable artifacts when the camera rotation leads to a significant deviation from the original view. In contrast, video generation methods like SEVA (Zhou et al. 2025b) and DimensionX (Sun et al. 2024) lack geometric consistency. Moreover, as these methods are not end-to-end frameworks, they cannot model the geometric interaction between humans and the scene, resulting in unrealistic compositions. In comparison, our method outperforms others in both geometric consistency and image quality. This demonstrates that our approach can provide camera controllable backgrounds for 4D human sequences within a fully automated synthesis pipeline.
4.3 Ablation study
We ablate key modules in Fig. 4. Replacing our 3D placement network with a naïve 2D depth-sampling baseline (rows 1–2) makes the avatar sink or float, whereas our method yields stable, scene-anchored poses. Substituting our joint post-reconstruction with VistaDream’s pipeline (rows 3–4) leaves black cavities beside the sofa and on the road, but our full model inpaints these gaps for a seamless scene. The qualitative differences are striking even under casual inspection and persist across a wide variety of indoor and outdoor scenes. A quantitative summary in the supplementary material further corroborates these visual findings, confirming that both 3D placement and view-fusion inpainting are critical for high-quality dynamic scene generation.
5 Conclusion and future work
AnimateScene converts a single scene image, human image, motion clip, and camera path into a controllable 4D video. It first matches the actor’s appearance to the scene lighting, then embeds a depth-guided 3D avatar, and finally inpaints view-dependent gaps for seamless composition. Extensive quantitative tests and a user study show AnimateScene surpasses state-of-the-art baselines in noise suppression, edge sharpness, structural consistency, detail, perceptual quality, and geometric coherence. Future work will extend to large-scale outdoor scenes with dynamic lighting and materials, support multi-actor interactions, enable real-time reconstruction and rendering, and incorporate language- and semantics-driven controls for high-level edits and narrative direction.